Commercially available sulphuric acid contains % 91 acid by mass and has a density of 1.83 g ML a ) calculate the molarity of the solution b ) volume of concentrated acid required to prepare 3.5 L of 0.50 M H2SO4

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
91 % acid by mass means 91 g acid in 100 g solution.
Density = 1.83 g/mL
Volume of solution = mass/density
  = 91/1.83
  = 49.73 mL

Molar mass sulpuric acid, H2SO4  = 2 x 1 + 32 + 4 x 16 = 98 g/mol
Number of moles of sulphuric acid = mass/molar mass
  = 91/98
  = 0.93

Molarity = number of moles/ volume of solution in litres
  = 0.93/0.0497
  = 18.71 M

Applying,
M1V1= M2V2
18.71 x V = 0.5 x 3.5
  V = 0.093 L
    = 93 mL

Thus, 93 ml of concentrated acid is required to prepare 3.5L of 0.50 M H2SO4
 
 
Regards
Arun (askIITians forum expert)